# Thermoelectric processes of quantum normal-superconductor interfaces

### v1

@@ Acknowledgements
We acknowledge funding from the Spanish Agencia Estatal de Investigación projects No. PID2022-142911NB-I00,  PID2020-117992GA-I00,  PID2023-150224NB-I00 and  CNS2022-135950, from the Spanish CM “Talento Program” project No. 2019-T1/IND-14088 and No. 2023-5A/IND-28927, and through the “María de Maeztu” Programme for Units of Excellence in R&D CEX2023-001316-M. L.A. acknowledges support from CONICET, Argentina as well as FonCyT, Argentina, through Grant No. PICT 2020-A-03661. A.B. acknowledges the MUR-PRIN 2022—Grant No. 2022B9P8LN-(PE3)-Project NEThEQS “Non-equilibrium coherent thermal effects in quantum systems” in PNRR Mission 4-Component 2-Investment 1.1 “Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)” funded by the European Union-Next Generation EU and CNR project QTHERMONANO. A.B. acknowledges discussions with A. Jordan, F. Taddei, M. Governale, F. Giazotto and the kind hospitality by the Institute of Quantum Studies at Chapman University where partially the work was done.

### v2

@@ Acknowledgements
We acknowledge funding from the Spanish Agencia Estatal de Investigación projects Nos. PID2022-142911NB-I00, PID2020-117992GA-I00, PID2023-150224NB-I00 and CNS2022-135950, from the Spanish CM “Talento Program” project No. 2019-T1/IND-14088 and No. 2023-5A/IND-28927, and through the “María de Maeztu” Programme for Units of Excellence in R&D CEX2023-001316-M. L.A. acknowledges support from CONICET, Argentina as well as FonCyT, Argentina, through Grant No. PICT 2020-A-03661. A.B. acknowledges the MUR-PRIN 2022—Grant No. 2022B9P8LN-(PE3)-Project NEThEQS “Non-equilibrium coherent thermal effects in quantum systems” in PNRR Mission 4-Component 2-Investment 1.1 “Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN)” funded by the European Union-Next Generation EU and CNR project QTHERMONANO. A.B. acknowledges discussions with A. Jordan, F. Taddei, M. Governale, F. Giazotto and the kind hospitality by the Institute of Quantum Studies at Chapman University where partially the work was done.